40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 16 23:25:07 2017, gautam, Update, SUS, MC1 SUS electronics investigation 20170116_231625.pngIMG_7175.JPGIMG_7174.JPG
    Reply  Tue Jan 17 21:29:52 2017, gautam, Update, SUS, MC1 SUS electronics investigation 
       Reply  Wed Jan 18 11:40:54 2017, gautam, Update, SUS, MC1 SUS electronics investigation MC1glitches.pngMC1coilGlitches.pngMC1CoilGlitchZoomed.png
          Reply  Wed Jan 18 14:23:47 2017, gautam, Update, SUS, MC1 SUS electronics investigation 
             Reply  Wed Jan 18 18:44:53 2017, gautam, Update, SUS, MC1 SUS electronics investigation MCsatBoxSwitch.png
                Reply  Thu Jan 19 08:25:12 2017, Steve, Update, SUS, MC1 SUS electronics investigation MC1_MC3_ITMY_ETMX_sensors.pngsensors_UL.png
                   Reply  Thu Jan 19 12:00:10 2017, gautam, Update, SUS, MC1 SUS electronics investigation SatBoxSwitchGlitch.png
                      Reply  Thu Jan 19 19:56:09 2017, rana, Update, SUS, MC1 SUS electronics investigation 
                         Reply  Fri Jan 20 11:16:30 2017, gautam, Update, SUS, MC1 SUS electronics investigation MCglitchesReturn.png
                            Reply  Mon Jan 23 10:24:01 2017, Steve, Update, SUS, MC1 SUS electronics investigation 3glitchingSUS.png
Message ID: 12742     Entry time: Fri Jan 20 11:16:30 2017     In reply to: 12741     Reply to this: 12745
Author: gautam 
Type: Update 
Category: SUS 
Subject: MC1 SUS electronics investigation 

Both suspensions have been relatively well behaved for the best part of the last two days, since I effected the Satellite Box swap. Today morning, I set about re-enabling the damping and locking the MC. Judging by the wall StripTool, it stayed locked for about 30 mins or so, after which the glitching returned.

Attached is a screenshot of the sensor signals from MC1 and MC3 (second trend), and also the highest band (>30Hz) BLRMS output for the same 10 channels (full data sampled at 16Hz). Note that MC1 and MC3 satellite boxes remain swapped. So the glitches now have migrated to the MC3 channels.

I need to think about whether this is just coincidence, or if me re-enabling the damping has something to do with the re-occurrence of the glitching...


Addendum 4.30pm: I've also re-aligned the Y arm. Its alignment has been stable over the last few hours, despite several mode cleaner lock losses in between, it recovers good IR transmission. The X arm has been re-aligned to green, but I can't get it locked to the IR - everytime I turn the LSC to ETMX on, there seems to be some large misalignment applied to it. c1iscaux was dead, I restarted it by keying the crate. I haven't had time to investigate the X arm locking in detail, I will continue to debug this.

ELOG V3.1.3-